OpenStudy (anonymous):

A conveyor belt carries supplies from the first floor to the second floor, which is 12 feet higher. The belt makes a 60° angle with the ground. How far do the supplies travel from one end of the conveyor belt to the other? Round your answer to the nearest foot.

OpenStudy (anonymous):

do you mean the length of the actual conveyor belt or the distance it travels on the x axis of a graph

OpenStudy (anonymous):

the length of the actual conveyer belt

OpenStudy (anonymous):



OpenStudy (anonymous):

yeah thats what i drew on my sheet

OpenStudy (anonymous):

tuffie

OpenStudy (anonymous):

\[\frac{12}{h}=\frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

\[h=\frac{24}{\sqrt{3}}\]

OpenStudy (anonymous):

good idea

OpenStudy (anonymous):

approx. 13.86 feet

OpenStudy (anonymous):

thanks!!

OpenStudy (anonymous):

hey satelite you agree or is mh math wrong